Lindenauer Corp. bought a machine on January 1, 2008 for $800,000. The machine had an expected life of 20 years and was expected to have a salvage value of $40,000. The company does not plan to dispose of the machine but does believe it may be impaired. On July 1, 2018, the company reviewed the potential of the machine and determined that its future net cash flows totaled $350,000 and its fair value was $230,000.

1. Calculate the book value of the Lindenauer’s machine immediately prior to the impairment test. (Hint: you will need to consider a partial year depreciation catch-up entry for the first half of 2018.
Answer #1

Cost of machine = (800,000-40,000)/20

Depreciation for the year = $ 38000

Accumulated depreciation from Year 2008 to 2017 = 10*38,000= $380,000.000

For 2018 half of the depreciation needs to be considered = 38,000/2= $19,000

Total Accumulated depreciation till June 2018 = $380,000+$19,000= $399,000.00

Book Value of the asset prior to impairment test = $401,000.00